OpenStudy (anonymous):

A wheel accelerates from rest for 2.0s with a constant acceleration of 3.0 rad/s^2. If the wheel has a radius of 0.50m, through what distance in m has a point on the rim moved?

OpenStudy (anonymous):

\[\theta= \omega_0 t +\frac{1}{2} \alpha t^2\] \[\theta= 0 t +\frac{3}{2} t^2\] t=2 \[\theta= \frac{3}{2} 4=6rad\] \[s=(\frac{\theta}{2\pi}) r=\frac{6}{2 \pi}0.5\]

OpenStudy (anonymous):

Thanks